OpenStudy (anonymous):

MEDAL GIVEN TO CORRECT ANSWER. A sled of mass 28 kg slides across the ground with a friction force of 200.3 N. What is k between the sled and the ground? A. 0.14 B. 0.73 C. 7.2 D. 1.2

OpenStudy (anonymous):

What are the Forces acting on the sled? The Force of Friction pointing opposite of the direction of movement, The Force of Weight pointing downwards, and the Force of Normal (Force of the surface on the object) pointing upwards. DRAW FREE BODY DIAGRAM! What is the equation for friction? Friction = (Normal)(Constant) Where can we get Normal from since we already know what Friction is? From the Free Body Diagram we see a Force of Normal pointing upwards How do we obtain a number value for that Normal? Since the sled isn't bobbing up and down, it doesn't have an acceleration in the y direction! Using Newton's First Law of Force Net = (Mass)(Acceleration) we find that with no acceleration the total Force in the y direction is zero. What are the total Forces in the y direction that make up the net Force we just solved for? (Force of Normal)+ (-Force of Weight) = Force Net We put a negative in front of the Force of Weight because it is pointing downwards and hinders the Force of Normal from pushing the sled too up. This equation is equal to zero. We find that the Force of Normal is equal to the Force of Weight Plug Normal into Friction equation and find the constant value of 0.73

OpenStudy (anonymous):

*What is the Equation of the Force of Weight? Weight = (mass)(gravity)

OpenStudy (anonymous):

Is the answer .73? I did the work and that's not what I got. @elsaj82

OpenStudy (anonymous):

Hmmmm... let me work it again F= (N)(k) N=W=(m)(g)= (28)(9.81)=274.68 200.3=(274.68)(k) k=.73 If you have anymore questions, fell free to ask!